Determine the molecular formula of a compound having a simplest formula of CH


and formula mass of 78.

1
Expert's answer
2022-06-08T01:42:04-0400

Solution:

Simplest formula mass of CH = Ar(C) + Ar(H) = 12 + 1 = 13 g mol−1

Molecular formula mass of compound = 78 g mol−1


Molecular formula mass / Simplest formula mass = n formula units/molecule

(78 g mol−1) / (13 g mol−1) = 6 formula units/molecule

 

Finally, derive the molecular formula of the compound from the simplest formula by multiplying each subscript by six: (CH)6 = C6H6

Therefore, the molecular formula of the compound is C6H6


Answer: The molecular formula of the compound is C6H6